OpenStudy (anonymous):

Find the direct variation equation of the graph through the points (0, 0) and (1, -2). Write in y=kx form.

OpenStudy (goformit100):

@Rufusmith

OpenStudy (anonymous):

y = -2x

OpenStudy (anonymous):

Ooh, okay! Gah before I put y = -1/2x :P Thank you(:

OpenStudy (anonymous):

you have to remember that y = mx +h, where: \[m = \frac{ \Delta y }{ \Delta x }\]

OpenStudy (anonymous):

i this case m = -2/1 and h = 0 So y = -2x

OpenStudy (anonymous):

Oohh, okay. My teacher didn't explain it that way:P

OpenStudy (anonymous):

xD